ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: Yang, Chung-Chun, The Hong Kong University of Sci. & Tech. Some recent results on factorization of meromorphic functions In this talk definitions and results on common right factor and common multiplier of two entire functions will be introduced. Specifically, the existence of a transcendental common right factor will be reported. Part of these results is joint work with T. W. Ng.